If sin (B) = -1/3 with B with B in Quadrant 3, find tan (B/2).


1
Expert's answer
2021-06-01T18:32:16-0400

If sin (B) = -1/3 and it is in 3 Quadrant where sinus turns negative. Then B = arcsin(-1/3)= arcsin(- 0.3333)= - 0.3398. Then tan(B/2)=tan(- 0.3398/2) =tan(- 0.1699) =- 0.002965 ≈ - 0.003.